# Modeling the Dynamics of Sub-Millisecond Electroadhesive Engagement and Release Times

### v1

@@ II-D Full Electromechanical Dynamics Model
$0=F_{N,base}+F_{ea}^{\prime}-F_{k}+F_{b}^{\prime}-m_{s}g$ (14)
where $F_{N,base}$ is the normal force from the base on the substrate and $m_{s}$ is the mass of the substrate. Since $F_{N,base}$ does not directly affect the dynamics, we leave it as is instead of modeling its contact mechanics and air gap.
@@
III Shear Force Capacity Results
III-A Experimental Setup and Testing Procedure
@@
Figure 5: Experimental test setup for measuring EA shear force as the load cell on a motorized stage pushes against the substrate.
@@
The EA pad was made using a 24 µm thick P(VDF-TrFE-CFE) film with 7 molCFE and a sputtered gold layer on one side (PolyK Technologies, LLC, USA). An interdigitated electrode pattern with dimensions in Table I was etched into the gold using a quasi-CW DPSS UV laser cutter (Series 3500, DPSS Lasers Inc., USA). Enameled copper wire was adhered to each electrode using silver conductive epoxy. As shown in Fig. 1(b), the interdigitated electrodes were designed to be wider than the substrate to mimic prior work on using EA clutches to manipulate small metallic objects [8], and to mitigate inconsistencies if the pin slipped sideways when reset between runs. The high voltage drive signal was powered by a 50 mA rated DC-DC converter (AHV12V500V50MAW, Analog Technologies, USA) and a high voltage shift register with a half-bridge output stage (HV507, Microchip Technology Inc., USA). We observed rise times (10% - 90%) of 8.3 µs and fall times of 5.2 µs when the power supply drove our EA pad. We measured an average power consumption of 1.43 mW for a 300 V, 1 kHz bipolar square wave drive signal.
@@
We used six brass substrates (Albion Alloys, England) with different dimensions shown in Table I. Screws were tightened through mounting points in the assembly to lightly press the dielectric into the substrate, which helped reduce film wrinkling under high shear stress. Pressure was evenly distributed along the sides using 3D printed brackets and a 1/16” closed-cell blended EPDM foam (McMaster-Carr Supply Company, USA). The foam was chosen for its low modulus (comparable to 5A durometer silicone) and low tackiness.
@@
The base that the substrate moves on was 3D printed from VeroWhitePlus photopolymer using a Strasys Objet24 V3. To measure the coefficient of friction, we took the ratio between tangential and normal force across 10 trials, measuring $\mu_{base}^{static}=0.281$ and $\mu_{base}^{kinetic}=0.173$ between the brass substrate and the base plate; and $\mu_{d}^{static}=0.188$ and $\mu_{d}^{kinetic}=0.154$ between the substrate and the dielectric film.
Figure 6: (a) Load cell force measured over the course of a shear force capacity test. The drive signal, shown for one electrode, was a 300 V, 10 Hz bipolar square wave. (b) Calculation procedure for the engagement time of this sample. The force profile’s rounded corner when the voltage rises is due to filtering, so the engagement time is measured when the force profile after the voltage rises linearly interpolates to the average force before the voltage rises (shown via dotted black lines). (c) Calculation procedure for the release time of this sample, measured as the time between when the voltage falls and when the load cell falls to within 10% of its final settling value. Using these metrics, for the sample shown $t_{engage}$ = 49.4 µs and $t_{release}=$ 1.49 ms.
Fig. 6 shows a sample experimental run. The stepper motor drives the substrate forward at 0.5 mm/s for 3 sec to establish a baseline kinetic friction force reading. We then turned on the voltage while continuing to drive the motor for 1.5 sec. Finally, the motor is stopped, the voltage is removed, and the load cell is allowed to settle for 2 sec. Load cell readings were filtered through a zero-phase 250 Hz low-pass filter to eliminate background noise, similar to prior work [6].
III-B Fitting Contact Mechanics Model
Figure 7: For a range of applied normal loads, we measured the capacitance and used it to estimate the air gap using Eq. 2. The average and standard deviation across 3 trials is shown.
@@ III-B Fitting Contact Mechanics Model
To fit the parameters $k$ and $\sigma_{d}$ in Eq. 8, we measured the capacitance of our EA pad against brass substrates of different widths across a range of applied normal forces, as shown in Fig. 7. The EA pad’s capacitance was read using a Keysight U1701B capacitance meter, after subtracting parasitic capacitance. The air gap was then estimated using Eq. 2. Fitting the air gap and normal force data to Eq. 8, our best fit parameters were $k=5.39\cdot 10^{14}$ N/m3.5 and $\sigma_{d}=2.80$ µm. This method is an accessible alternative to traditional surface profilometry for characterizing EA dielectric surfaces [44].
III-C Shear Force Capacity Measurements and Modeling
@@
Although maximizing shear force capacity is not a primary objective of this paper, understanding how different parameters affect our clutch’s force capacity can help validate whether our model is physically sound. Fig. 8 shows the measured and predicted (Eq. 16) shear force capacities at different voltages and bipolar square wave AC drive frequencies.
@@
When applying our model, we found that our analytical EA force equation was slightly inconsistent from what was needed to produce the measured shear capacities, so we fitted a multiplier $\lambda^{fit}_{ea}(V)$ for $F_{ea}$ to each voltage’s curve in Fig. 8:
@@
We found that the best fit values followed a roughly linear relationship $\lambda^{fit}_{ea}(V)=2.40-0.0058V$ ($R^{2}=0.98$). The average value of $\lambda^{fit}_{ea}$ over all samples was $1.08$ $(\sigma=0.33)$. Because the average $\lambda^{fit}_{ea}$ was about 1 and because the model accurately predicted the shear pressure capacity’s decay at high frequencies, we take these results as confirmation that our electromechanical model is physically sound, except for what we hypothesize are unmodeled parasitics.
@@
The -3 dB frequency for all voltages is experimentally over 10 kHz, and is predicted by our model to be about 13.3 kHz. For comparison, Chen et al. [19] measured a -3 dB frequency of about 150 Hz for both square and sine AC drive signals using a glass substrate, Zhang et al. [8] measured a -3 dB frequency of about 130 Hz for sine AC drive signals against a brass metal substrate, and Shultz et al. [10] measured a -3 dB frequency of around 75 Hz for sine AC drive signals using a human fingertip as the substrate. We predict the difference is due largely to the improved dynamics of our drive circuitry and our use of a metal substrate.
@@
Figure 8: Shear force capacity measured for a 2 mm wide substrate at different voltages and drive frequencies, averaged over at least 5 trials per condition.
III-D Force Slip vs. Frequency
Figure 9: Max slip percentage measured for a 2 mm wide substrate at different voltages and drive frequencies, averaged over at least 5 trials per condition.
@@
As shown in the sample force profile in Fig. 6(a), EA clutches lose a significant amount of their load-carrying capacity after a slip, with similar clutches in the literature often losing 50-70% of their measured force and requiring tens of milliseconds before the clutch re-engages [6, 24]. This failure mode is a major deterrent to using EA clutches in safety-critical robots. Fig. 9 shows how the worst slip in each of our test runs compares across different voltages and frequencies, measured as the percentage of force lost after the slip. We found that the slip percentage decreases drastically at higher frequencies, with 250 V seeing a 4.7$\times$ reduction on average between the 10 Hz and 5 kHz drive frequencies. Our numerical values cannot be compared directly with prior work, since the slip percentage also varies with load cell dynamics, but prior work has primarily used DC up to 10 Hz drive signals. Our hypothesis, which we explore further in Sec. IV, is that higher drive frequencies can enable faster actuation times.
IV Engagement and Release Time Results
IV-A Engagement Time Testing and Simulation Methods
To measure the engagement time, we adapted a method from Diller et al. [6], shown in Fig. 6(b). We linearly interpolated the load cell readings after the voltage is enabled down to the average force over the 10 ms before the voltage is enabled. To minimize the impact of noise, we found linear interpolations for every data interval starting from 0.5 ms after the voltage is enabled and ending between 2-20 ms after the voltage is enabled. After filtering to only fits with $R^{2}>0.8$, we chose the smallest positive interpolated intersection to be the engagement time. For some tests, there would be no positive values due to noise; in these cases, we rounded the run’s measured engagement time up to 0.
IV-B Engagement Time Results and Discussion
Fig. 10 shows the engagement time across a total of 311 test runs over different drive voltages and frequencies, substrate widths, and preload forces, showing in general that the EA clutch can engage within several tens of microseconds. This matches our hypothesis when modeling that EA actuators can be designed to engage about as quickly as it takes to polarize their dielectric (for comparison, our power supply had a 10% - 90% rise time of 8.3 µs, and our dielectric had a polarization relaxation time of $\tau=2.82$ µs). Across all 311 test runs, the average engagement time was 4.78 µs ($\sigma$ = 7.07 µs), and 89.7% (279/311) of our measured engagement times were $<$15 µs.
We observed significant run-to-run variance, which is not uncommon in the literature given the stochastic nature of the interaction between dielectric asperities and the substrate [6, 8]. We also attribute this noise in part to our ADC. According to its datasheet, our ADC has 19.3 effective bits of resolution at our sample rate and gain, which allows our setup to detect force changes of about 125 µN. At our experimental setup’s movement speed (0.5 mm/s) and load cell spring constant ($k_{lc}=18.0$ N/mm), this corresponds to a delay of about 13.9 µs before we can expect to see a measurable change in force after the voltage engages. This aligns well with the range of standard deviations in our experimental results.
Comparing results for our $1.11-3.33$ cm2 EA-ICE clutches to prior literature, Hinchet and Shea [24] achieved engagement times of 5 ms for a 3 cm2 parallel plate EA clutch, while Zhang et al. [8] achieved an engagement time of 6.7 ms for an 0.96 cm2 EA-ICE clutch. In 1957, Fitch [25] achieved an engagement time of 150 µs for a 38 cm2 parallel plate rotary clutch using silicone oil as the dielectric and very precisely machined electrodes, which we believe helped reduce their interfacial surface roughness. While we are not able to directly apply our model to prior work, given our lack of contact models for their dielectric-substrate interfaces, we explore several hypotheses for why similarly sized EA clutches might see much larger engagement times in Sec. V.
Our model’s predicted engagement times generally followed the trends of our experimental data. In Fig. 10(a), linearly interpolating the experimental data gives a similar slope -0.0013 µs/V compared to our model’s predicted slope -0.0012 µs/V. Both Diller et al. [6] and Zhang et al. [8] similarly observed that higher voltages corresponded to lower engagement times. In Fig. 10(b), linearly interpolating the experimental data gives a slope 1.15 µs/mm compared to our model’s predicted slope 0.98 µs/mm. Notably, although the EA normal force is linearly proportional to the substrate width, we see a positive correlation between engagement time and substrate width instead of a negative one because the damping force scales as the cube of substrate width. In Fig. 10(c), as predicted by our model, we did not observe a consistent trend between engagement time and drive frequency because our measured engagement times were less than the period of any of our tested drive signals. In Fig. 10(d), our model predicted a small increase in engagement time for zero preload normal force, but it quickly steadied for higher preload forces, and we observed no consistent trend experimentally.
IV-C Release Time Testing and Simulation Methods
@@
As shown in Fig. 6(c) and consistent with previous work [6, 24], we measured the release time for our EA clutches as the time between when the voltage begins to fall and when the load cell falls 90% of the way to its final steady-state value.
@@
$m_{lc}\ddot{x}_{lc}=-k_{lc}x-b_{lc}\dot{x}_{lc}-F_{shear}\textrm{sign}(\dot{x}_{lc})$ (19)
where $F_{shear}$, given by Eq. 15, is calculated using static friction coefficients when $\dot{x}=0$ and kinetic friction coefficients otherwise. We measured the load cell’s spring constant ($k_{lc}=18.0$ N/mm), damping constant ($b_{lc}=0.43\frac{N\cdot s}{m}$), resonance frequency ($\omega_{lc}=643.4$ Hz), and effective mass ($m_{lc}=1.10$ g) by recording its response to step and ramp inputs to the load cell’s position by the stepper motor. While the load cell dynamics were important to include in the model to fit the data, the load cell’s quarter-period $\omega_{lc}/4=388.6$ µs is small enough that we can expect the clutch’s electromechanical dynamics to be the primary contributor to the measured release time.
For each experimental run, we estimated $F_{preload}$ by measuring the kinetic friction on the substrate during the first 3 sec before the voltage is enabled and applying Eq. 17. We also measured the ratio $\frac{F_{shear}(t_{r})}{F_{shear,max}}$ between the load cell’s applied force right before the voltage was released and the load cell’s maximum measured force. Using $F_{preload}$, we solved the coupled equations of motion Eqs. 13 and 14 to calculate $T_{air}(t_{r})$ at the moment of release. We then estimated the shear force capacity $F_{shear,max}^{predicted}$ of the EA pad using Eq. 16. We estimated the initial position of the load cell as:
@@
IV-D Release Time Results and Discussion
Fig. 11 shows release time results for different geometric and operating parameters. The best release time was 874.6 µs ($\sigma=$ 31.2 µs) for a 2 mm wide substrate, a low preload force (averaged over test runs with $F_{preload}<0.25$ N), and a 225 V, 1 kHz drive signal. Across all 311 runs in our dataset, the average release time was 1.13 ms ($\sigma=$ 0.36 ms), and 39.5% of the runs had release times $<1$ ms.
Our release times are much faster than those in prior work. Hinchet and Shea [24] reported release times around 15 ms for a 3 cm2 parallel plate clutch, while Diller et al. [45] measured 23.6 ms release times for a 80 cm2 parallel plate clutch. The best EA-ICE clutch result we found was by Cao et al. [33], who measured release times of 100 ms for a 24.6 cm2 clutch.
In Fig. 11(a), linearly interpolating the experimental data gives a slope -0.72 µs/V compared to our model’s predicted slope -0.25 µs/V. In Fig. 11(b), linearly interpolating the experimental data gives a similar slope 143.5 µs/mm compared to our model’s predicted slope 142.3 µs/mm. In prior work, Diller et al. [6] similarly observed that EA release times decrease with voltage and increase with substrate width.
In Fig. 11(c), we observed a local minima for release times, with 2 kHz drive frequencies causing on average 44.2% and 10.5% lower release times than 1 Hz and 10 kHz drive frequencies respectively. We did not find a model that captures the low-frequency trend. Our original hypothesis was that our drive circuit might be leaking charge into the EA clutch, but Fig. 8 shows our clutch’s shear force capacity did not increase at low frequencies. For drive frequencies above 2 kHz, however, our model did predict our experimental trend towards higher release times. At high frequencies, the EA force capacity decays, increasing the impact of slower mechanical dynamics in the release time. Notably, our experimental trends for release time line up very well with the trends in slip percentage from Fig. 9, where we saw the percentage of force lost after a slip decrease significantly as the drive frequency increased from 10 Hz to 2 kHz but increase slightly as the drive frequency increased to 10 kHz. We hypothesize that because our EA clutches can release faster around 2 kHz, at those drive frequencies it can also more quickly dissipate the interfacial static charges that accumulate after a slip and recover faster.

### v2

@@ 2.4 Full Electromechanical Dynamics Model
$0=F_{N,base}+F_{ea}^{\prime}-F_{k}-F_{b}^{\prime}-m_{s}g$ (14)
where $F_{N,base}$ is the normal force from the base on the substrate and $m_{s}$ is the mass of the substrate. Since $F_{N,base}$ does not directly affect the dynamics, we leave it as is instead of modeling its contact mechanics and air gap. The system of equations given by Eqs. 13 and 14 can be solved as an initial value problem with the state $[T_{air},\dot{T}_{air}]$.
@@
3 Shear Force Capacity Results
3.1 Experimental Setup and Testing Procedure
To test our model, we built the experimental test setup in Fig. 5(a) to measure the shear force capacity, engagement time, and release time of an EA pad to a metallic substrate across a variety of geometric and operating parameters. Similar to prior work on both parallel plate EA clutches [7, 20] and EA-ICE clutches [9, 45, 62], we created a tensile testing machine by mounting a 500g load cell (UXCell, China) on a motorized linear slide. The stepper motor moving the slide block was driven at 1/256 microstepping, and the load cell was sampled at 38.4 kHz (ADS1263, Texas Instruments, USA).
@@ 3.1 Experimental Setup and Testing Procedure
The EA pad was made using a 24 µm thick P(VDF-TrFE-CFE) film with 7 molCFE and a sputtered gold layer on one side (PolyK Technologies, LLC, USA). An interdigitated electrode pattern with dimensions in Table 1 was etched into the gold using a UV laser cutter (Series 3500, DPSS Lasers Inc., USA). Enameled copper wire was adhered to each electrode using silver conductive epoxy. As shown in Fig. 1(b), the interdigitated electrodes were designed to be wider than the substrate to mimic prior work on using EA clutches to manipulate small metallic objects [9], and to mitigate inconsistencies if the pin slipped sideways when reset between runs. The high voltage drive signal was powered by a 50 mA rated DC-DC converter (AHV12V500V50MAW, Analog Technologies, USA) and a high voltage shift register with a half-bridge output stage (HV507, Microchip Technology Inc., USA).
@@
We used six brass substrates (Albion Alloys, England) with different dimensions shown in Table 1. Screws were tightened through mounting points in the assembly to lightly press the dielectric into the substrate, which helped reduce film wrinkling under high shear stress. Pressure was evenly distributed along the sides using 3D printed brackets and a 1/16” closed-cell blended EPDM foam (McMaster-Carr Supply Company, USA). The foam was chosen for its low modulus (comparable to 5A durometer silicone) and low tackiness. This setup constrains the brass substrate to move primarily in the $x$ direction and constrains the dielectric film to move primarily in the $z$ direction, in line with our model assumptions in Sec. 2. The boundary constraints also apply a small preload force $F_{preload}<0.25$ N to our dielectric film, which is among the lower range of preload forces used in prior work [9, 11, 19, 34, 61]. Almost all prior works in Fig. 1(b) apply a preload force to ensure conformality and prevent dielectric film wrinkling. Although smaller preload forces tend to be better for dynamics measurements to avoid additional sources of noise, in Sec. 4, we explore adding extra clamping pressure to test the effect of higher preload forces within the range seen in prior work.
@@
The base that the substrate moves on was 3D printed with a glossy finish from VeroWhitePlus photopolymer using a Strasys Objet24 V3. To measure the coefficient of friction, we took the ratio between tangential and normal force across 10 trials, measuring $\mu_{base}^{static}=0.281$ and $\mu_{base}^{kinetic}=0.173$ between the brass substrate and the base plate; and $\mu_{d}^{static}=0.188$ and $\mu_{d}^{kinetic}=0.154$ between the substrate and the dielectric film. Additional testing setup implementation details are included in Supplementary Materials, Sec. S4.
[…]
Fig. 5(c) shows a sample experimental run. We include the raw data for 10 additional experimental runs in Supplementary Materials, Sec. S5 and a video of a run in Supplementary Video 1. The stepper motor drives the substrate forward at 0.5 mm/s for 3 seconds to establish a baseline kinetic friction force reading. We then turned on the voltage while continuing to drive the motor for 1.5 seconds. Finally, the motor is stopped, the voltage is removed, and the load cell is allowed to settle for 2 seconds. We chose to push the rigid substrate instead of pulling on our soft dielectric film to avoid plastic deformation, in line with similar guidance by the ISO 8295 standard on measuring the frictional properties of thin films [63]. A timing diagram of the actuation and sensing signals is shown in Supplementary Materials, Sec. S4.1. The ADC is synced using the Teensy microcontroller to always sample the load cell and high voltage drive signal after the stepper motor settles its position, and the first reading is timed to occur about 4 µs after the high voltage first turns on.
3.2 Fitting Contact Mechanics Model
Figure 6: (a) Experimental test setup for measuring the EA pad capacitance over a range of applied normal forces. Given the measured capacitance values shown in (b), Eq. 2 was used to estimate the air gaps as shown in (c). The average and standard deviation across 3 trials are shown. The contact mechanics model in Eq. 8 was least squares fit to the curves in (c). Using these fitted parameters, the model’s predicted capacitances across the normal force range are also plotted using dashed lines in (b).
@@ 3.2 Fitting Contact Mechanics Model
To fit the parameters $k$ and $\sigma_{d}$ in Eq. 8, we measured the capacitance of our EA pad against brass substrates of different widths across a range of applied normal forces, as shown in Fig. 6. The EA pad’s capacitance was read using a Keysight U1701B capacitance meter, after subtracting parasitic capacitance. The air gap was then estimated using Eq. 2. Fitting the air gap and normal force data in Fig. 6(c) to Eq. 8, our best fit parameters were $k=5.39\cdot 10^{14}$ N m-3.5 and $\sigma_{d}=2.80$ µm. This calibration procedure is inspired by prior work on capacitive contact profilometry [53, 54, 55, 56] and presents an accessible alternative to traditional optical or stylus-based surface profilometry for characterizing EA dielectric surfaces [60, 62].
As a technical note, we also tried fitting our capacitance data in Fig. 6(b) to the surface roughness-averaged capacitance using a similar integral formulation as Eq. 12. However, our best fit parameters were almost identical to the ones obtained by our parallel plate capacitance model in Eq. 1 and Eq. 2. This aligns with expectations based on prior work applying capacitive contact profilometry to surfaces with micrometer-scale surface roughnesses [53, 54, 55, 56]. It also matches the methods used by prior EA studies to estimate their clutches’ average air gap [11, 34, 57, 58].
3.3 Shear Force Capacity Measurements and Modeling
@@
Although maximizing shear force capacity is not a primary objective of this paper, understanding how different parameters affect our clutch’s force capacity can help validate whether our model is physically sound. Fig. 7(a) shows the measured and predicted (Eq. 16) shear force capacities at different voltages and bipolar square wave AC drive frequencies. The order of voltages and frequencies tested was randomized.
@@
When applying our model, we found that our analytical EA force equation was slightly inconsistent from what was needed to produce the measured shear capacities, so we fitted a multiplier $\lambda^{fit}_{ea}(V)$ for $F_{ea}$ to each voltage’s curve in Fig. 7(a):
@@
We found that the best fit values followed a roughly linear relationship $\lambda^{fit}_{ea}(V)=2.40-0.0058V$ ($R^{2}=0.98$). The average value of $\lambda^{fit}_{ea}$ over all samples was 1.081.081.08 $(\sigma=0.33)$. For context, although fitted multipliers are almost universal in prior work to account for hard-to-model physical and electrical non-idealities [61], they tend to have a wide range of magnitudes ranging anywhere from $10^{-4}$ [64] to 10 [7]. Fitted voltage-dependent multipliers have seen particular success in recent studies [49, 65, 66, 67], drawing inspiration from the fact that many parasitics and dielectric behaviors like electrostriction are voltage-dependent. Because our average $\lambda^{fit}_{ea}$ was about 1 and because the model accurately predicted the shear pressure capacity’s decay at high frequencies, we take these results as confirmation that our electromechanical model is, for the most part, physically sound. We leave in-depth modeling of these voltage-dependent behaviors to future work, and for the purposes of this study we proceed with the fitted multiplier $\lambda_{ea}^{fit}(V)$.
@@
The -3 dB bandwidth for all voltages is experimentally over 10 kHz, and is predicted by our model to be about 13.3 kHz. For comparison, Chen et al. [18] measured a -3 dB frequency of about 150 Hz for both square and sine AC drive signals using a glass substrate, Zhang et al. [9] measured a -3 dB frequency of about 130 Hz for sine AC drive signals against a brass metal substrate, and Shultz et al. [11] measured a -3 dB frequency of around 75 Hz for sine AC drive signals using a human fingertip as the substrate. We predict the difference is due largely to the improved dynamics of our drive circuitry and our use of a metal substrate. Additional comparisons between our experimental results and prior literature are provided in Supplementary Materials, Sec. S6.
@@
Figure 7: (a) Shear force capacity and (b) max slip percentage measured for a 2 mm wide substrate at different voltages and drive frequencies, averaged over at least 5 trials per condition.
3.4 Force Slip vs. Frequency
@@
As shown in the sample force profile in Fig. 5(c), EA clutches lose a significant amount of their load-carrying capacity after a slip, with similar clutches in the literature often losing 50-70% of their measured force and requiring tens of milliseconds before the clutch re-engages [7, 20, 42]. This failure mode is a major deterrent to using EA clutches in safety-critical applications. Fig. 7(b) shows how the worst slip in each of our test runs compares across different voltages and frequencies, measured as the percentage of force lost after the slip. We found that the slip percentage decreases drastically at higher frequencies, with 250 V seeing a 4.7$\times$ reduction on average between the 10 Hz and 5 kHz drive frequencies. Our numerical values cannot be compared directly with prior work, since the slip percentage also varies with load cell dynamics, but prior work has primarily used DC or low-frequency ($\leq$10 Hz) AC drive signals. Our hypothesis, which we explore further in Sec. 4, is that higher drive frequencies should reduce space charge buildup in the dielectric, enabling smoother release and faster re-engagement.
4 Engagement and Release Time Results
Both faster engagement times and faster release times are important to improving the effective bandwidth of EA clutches. We focus first on the release times of our EA-ICE clutches because, as shown in Fig. 1(c), they tend to be much larger than the engagement times and thus a much larger component of EA bandwidth.
4.1 Release Time Testing and Simulation Methods
@@
As shown in Fig. 5(e) and consistent with previous work [7, 20], we measured the release time for our EA clutches as the time between when the voltage begins to fall and when the load cell falls 90% of the way to its final steady-state value.
@@
$m_{lc}\ddot{x}_{lc}=-k_{lc}x_{lc}-b_{lc}\dot{x}_{lc}-F_{shear}\textrm{sign}(\dot{x}_{lc})$ (19)
where $F_{shear}$, given by Eq. 15, is calculated using static friction coefficients when $\dot{x}_{lc}=0$ and kinetic friction coefficients otherwise. We measured the load cell’s spring constant ($k_{lc}=18.0$ N mm-1), damping constant ($b_{lc}=0.43$ kg s-1), resonance frequency ($\omega_{lc}=643.4$ Hz), and effective mass ($m_{lc}=1.10$ g) by recording its response to step and ramp inputs to the load cell’s position by the stepper motor. While the load cell dynamics were important to include in the model to fit the data, the load cell’s quarter-period $\omega_{lc}/4=388.6$ µs is small enough that we can expect the clutch’s electromechanical dynamics to be the primary contributor to the measured release time. Additional discussion motivating this conclusion is presented in Supplementary Materials, Sec. S7.
For each experimental run, we estimated $F_{preload}$ by measuring the kinetic friction on the substrate during the first 3 sec before the voltage is enabled and applying Eq. 17. We also measured the ratio $\frac{F_{shear}(t_{r})}{F_{shear,max}}$ between the load cell’s applied force right before the voltage was released and the load cell’s maximum measured force. Using $F_{preload}$, we solved the coupled equations of motion Eqs. 13 and 14 to calculate $T_{air}(t_{r})$ at the moment of release. We then estimated the shear force capacity $F_{shear,max}^{predicted}$ of the EA pad using Eq. 16. We estimated the load cell’s initial position as:
@@
4.2 Release Time Results and Discussion
Fig. 8 shows the release times of our EA-ICE clutches across a total of 311 test runs over different drive voltages and frequencies, substrate widths, and preload forces. The best release time was 874.6 µs ($\sigma=$ 31.2 µs) for a 2 mm wide substrate, a low preload force (averaged over test runs with $F_{preload}<0.25$ N), and a 225 V, 1 kHz drive signal. Across all 311 runs in our dataset, the average release time was 1.13 ms ($\sigma=$ 0.36 ms), and 39.5% of the runs had release times <1<1< 1 ms.
Our $1$ 3.33 cm2 EA-ICE clutches have much faster release times than other centimeter-scale EA clutches are reported to have in prior literature. Hinchet and Shea [20] reported release times around 15 ms for a 3 cm2 parallel plate clutch, while Diller et al. [68] measured 23.6 ms release times for a 80 cm2 parallel plate clutch. The best EA-ICE clutch result we found was by Cao et al. [29], who measured release times of 100 ms for a 24.6 cm2 clutch adhering to a cardboard substrate. Our results fall behind the release times of sub-millimeter-scale EA clutches like MEMS gap closing actuators and relays, which can take advantage of their low mass and smaller relative damping forces to achieve microsecond-scale release times [36, 69, 70].
In Fig. 8(a), linearly interpolating the experimental data gives a slope -0.72 µs V-1 compared to our model’s predicted slope -0.25 µs V-1. 300 V drive voltages caused, on average, 20.7% smaller release times than 125 V drive voltages did. In Fig. 8(b), linearly interpolating the experimental data gives a similar slope 143.5 µs mm-1 compared to our model’s predicted slope 142.3 µs mm-1. 2 mm wide substrates had 83.8% smaller release times on average than 6 mm wide substrates did. In prior work, Diller et al. [7] similarly observed that EA release times decrease with voltage and increase with substrate width.
In Fig. 8(c), we observed a local minima for release times, with 2 kHz drive frequencies causing on average 44.2% and 10.5% lower release times than 1 Hz and 10 kHz drive frequencies respectively. We did not find a model that captures the low-frequency trend. Our original hypothesis was that our drive circuit might be leaking charge into the EA clutch, but Fig. 7 shows our clutch’s shear force capacity did not increase at low frequencies. For drive frequencies above 2 kHz, however, our model did predict our experimental trend towards higher release times. At high frequencies, the EA force capacity decays, increasing the impact of slower mechanical dynamics in the release time. Notably, our experimental trends for release time line up very well with the trends in slip percentage from Fig. 7(b), where we saw the percentage of force lost after a slip decrease significantly as the drive frequency increased from 10 Hz to 2 kHz but increase slightly as the drive frequency increased to 10 kHz. We hypothesize that because our EA clutches can release faster around 2 kHz, at those drive frequencies it can also more quickly dissipate the interfacial static charges that accumulate after a slip and recover faster.